OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2). A.1/5 B.5 C.-1/5 D.-5

OpenStudy (anonymous):

y1-y2/ x1-x2

OpenStudy (zzr0ck3r):

(-2-8)/(6-4) = ?

OpenStudy (anonymous):

-2-8/ 6-4= -10/2

OpenStudy (anonymous):

-5/1